# Regularized Adjusted Plus-Minus Models for Evaluating and Scouting Football (Soccer) Players using Possession Sequences

### v1

@@ 5 Discussion
In this paper, we present a novel framework for evaluating the strength of football players. The idea of our approach is based on top-down approaches such as plus-minus (PM) models, however, we provide two novel adaptions. First, the setup relies on possession sequences, which are sequences of consecutive on-ball actions of a team. Thus, in contrast to classical PM approaches, much more granular segments of the game are used for deriving a rating, which allows for evaluating on-ball and off-ball contributions of players to the game. Second, we incorporate four different penalization structures into the regression setup. As a start, we follow the classical approach and estimate player strengths by using a ridge penalty term. This is similar to existing PM models, however analyzing possessions allows for interesting novel insights. Next, the natural grouping of players into position groups (forward, midfielder, defender, goalkeeper) is exploited via the group lasso penalty, which allows for ridge shrinkage between groups instead of all players jointly. The third penalization structure employed, the exclusive lasso, again utilizes the group structure but uses a lasso type of penalty within these, such that strong (resp. weak) players are selected, whereas average players are set to zero. Such group penalties have not been considered in the literature so far, even though it is well-known that the position of players is essential when evaluating them. As a final regularization structure, we apply the generalized lasso penalty. It implicitly tries to group players of equal strength together. This type of penalty has not yet been used for analyzing football players either and has in general not been studied very thoroughly in the context of binomial logistic regression. Therefore, we also outline a scheme for implementing such a constraint using a conic optimization approach.
The rating methods are analyzed in detail in Section 4. On the one hand, we provide comparison and football-specific implications of the methods and on the other hand, we try to objectively evaluate the validity of the rating approaches. In general, there are similarities in all the approaches. The on-ball contributions are dominated by offensive players (forwards and midfielders), while defensive players score better on the off-ball contributions. From a domain-specific viewpoint, this result is expected as the measurement for evaluating possession is whether they result in a goal or not. It is reassuring that all the methods act similarly and these results suggest that they are meaningful and informative. The distinction between active contributions and off-ball contributions can be leveraged in two ways. While it may be more reasonable to evaluate offensive players on their direct involvement contribution and defensive players on their indirect involvement, the other way around may be of interest to coaches or scouts. Considering defensive players who perform well in on-ball contributions allows for the identification of more offensively oriented defenders, which might be desirable for specific styles of play. Similarly, offensive players that demonstrate high off-ball value may have more impact on the game than their active contributions (as measured e.g. through goals, assists, etc.) are disclosing. In terms of the validity checks conducted in this study, the regularized regression models clearly and significantly outperform baseline as well as reference models. The best-performing ones are the group lasso closely followed by the ridge regression. These models are at large quite similar, however, it seems that there is a slight advantage of applying the ridge shrinkage group-wise, as is done by the group lasso. The exclusive lasso also performs fairly well in terms of validity. Moreover, it offers interesting insights that may especially benefit scouting decisions through its ability to select outstanding players of a predefined group. The generalized lasso penalty performs worst among the regularization methods, however it offers value in the interpretation of the results. The implicit grouping of players into equal-strength groups indicates that distinctions between average players have to be conducted with care.
In more general terms, while our results provide interesting and novel insights into the evaluation and the scouting of players our framework also offers the possibility to enhance plus-minus rating models via a new way of segmenting a football match as well as the usage of a variety of penalization structures. Kharrat et al. (2020) e.g. mention some pitfalls that have to be considered when employing their PM approach to evaluate players. With the possessions data at hand and grouping of players into position groups some of the pitfalls mentioned in their paper can be avoided.
@@
Furthermore, the flexibility of our approach combined with the ease of adapting it to existing procedures yields several possible future directions to boost football player performance evaluation. For one, it might be interesting to simply adapt PM models to allow for the presented regularization schemes. Also incorporating adjustments for home advantage, manpower, league strength or relevancy of performance (cf. Kharrat et al., 2020; Pantuso and Hvattum, 2021) is straightforward. However, this becomes more relevant when a larger amount of data is available. Another interesting direction is to adapt the target variable and allow for some sort of hybrid approach. One possibility is to follow Kharrat et al. (2020) and use xG values as target variable, however, any kind of bottom-up rating that assigns a value at the end of the possession may be used. One could even think of ordinal-type target variables, however in this case the application of the more elaborated penalization structures needs to be developed further. Adapting the target may be helpful in discriminating between players, as considering goals as outcome variables leads to a sparse and imbalanced setup. Similarly, since analyzing possessions provides more detailed insights into the game than analyzing segments as defined in PM models, it is possible to adjust the outcome variable to allow for the valuation of defensive actions such as the ability of a team to stop dangerous possessions. Finally, future endeavors may explore and compare different setups for incorporating player information. In this work, we introduced the idea of taking into account on-ball as well as off-ball actions into a top-down framework. Whereas we focus on modeling the on-field contribution additively, it might be an interesting avenue to account for the off-ball involvement by separating offensive and defensive participations on the field. For instance, a player like Lionel Messi, while without a doubt positively influencing an offensive possession simply by being a threat on the field, has arguably only little effect with regards to defensive off-ball contribution.
@@
In summary, we believe that the presented framework allows for novel perspectives and enhancements in the field of player evaluation and scouting. The results are not only from a domain-specific but also from an objective perspective useful, and thus the stakeholders involved into the sports of football such as clubs, betting companies, media as well as fans may benefit from our evaluation methods.
Appendix A Debiased machine learning models for rating players
Another approach for deriving a rating that differs from the regularized regression is presented in Bajons (2023). In this work, a debiased machine learning (DML) model for studying player strengths is introduced. The main advantage of this model over the regularized approaches is that resulting player strengths may be interpreted in a certain way, that is to some extent more intuitive than the regularized strength estimates.
A.1 The DML model and the PCV rating
Consider the following partially linear model
$Y=D\theta_{0}+g_{0}(X)+\epsilon,$ $\quad\mathbb{E}[\epsilon|D,X]=0,$ (18)
$D=m_{0}(X)+\nu,$ $\quad\mathbb{E}[\nu|X]=0,$
where $Y$ represents the outcome variable, $D$ is the treatment variable, which is potentially highly correlated with $X$, $X$ are confounders, $\theta_{0}$ is the parameter of interest, $\epsilon$ and $\nu$ are stochastic error terms. Note that $g_{0}$ and $m_{0}$ are (potentially non-linear) nuisance functions, i.e. we need to estimate them, however, this is not our primary interest. In our setup, $Y$ is the outcome variable of whether the possession resulted in a goal or not. The variable $D$ is the involvement variable of a specific player, that we want to estimate, and $X$ is the feature matrix as described in the previous section, such that $X_{mod}=(X,D)$. As $Y$ is a binary variable, the above model specification is clearly only an approximation to the true underlying data-generating process, nevertheless, it allows for nice interpretation. The estimate for $\theta_{0}$ in this case can be seen as the level shift in the probability of scoring a goal from a possession when the player corresponding to $D$ is involved in the possession.
In order to estimate the parameter $\theta_{0}$, Chernozhukov et al. (2018) propose a debiased machine learning approach, which provides means for decreasing the regularization and overfitting bias in the estimation procedure. Their approach is based on the idea of partialling out the effect of the confounders $X$ from equation (18), i.e. rewriting it in the form
$W=\nu\theta_{0}+\epsilon,$ $\quad\mathbb{E}[\epsilon|D,X]=0,$ (19)
$W=Y-\ell_{0}(X),$ $\quad\ell_{0}(X)=\mathbb{E}[Y|X]=m_{0}(X)\theta_{0}+g_{0}(X),$
$\nu=D-m_{0}(X),$ $\quad m_{0}(X)=\mathbb{E}[D|X].$
The estimation of the treatment effect $\theta_{0}$ thus amounts to executing 2 steps:
  1. 1.
    Estimate $m_{0}$ and $\ell_{0}$ using machine learning techniques of choice, i.e. predict $D$ and $Y$ using $X$.
  2. 2.
    Estimate $\theta_{0}$ by regressing the residuals $\hat{W}=Y-\hat{\ell}_{0}(X)$ onto the residual of $\hat{\nu}=D-\hat{m}_{0}(X)$.
Combining these two steps with a suitable cross-validation strategy can be shown to reduce regularization bias as well as overfitting bias, thus providing to some extent better estimates than using a direct approach for estimation of $\theta_{0}$. The above two-step procedure can be used for estimating the effect $\theta_{0}$, i.e. the procedure is repeated such that for every player $i$ in the dataset an estimate $\hat{\theta}_{0,i}$ is obtained. The resulting estimate is nicely interpretable as the level shift in the probability of scoring when player $i$ is involved. However, the raw estimate is still subject to problems that arise from the imbalance in the dataset. The effect for players with little involvement is overestimated, while the effect for players who highly participate in the game is underestimated. This is because it is more difficult to maintain a level of effectiveness when a player is a key player and receiving a lot of passes and thus exhibiting a high amount of involvements. Pure strikers for example mostly are only involved at the end of a play and do not have to make risky passes but rather convert the chances they have. Conversely, midfielders and creative players receive the ball far more often during the game. Furthermore, they have to take more risks by passing the ball through defenses or dribbling through a wall of defenders. Thus similarly estimated values of $\theta_{0}$ might hide the fact, that one player has a lot of involvements in a lot more build-up possessions than the other player and thus is much more important to the game and the team. In order to account for this fact and value players with respect to not only quality (respectively effectiveness) but also quantity, a transformation of the $\hat{\theta}_{0}$ value is proposed, which takes into account the amount of actions a player has. The metric for rating player $i$ termed possession contribution value (PCV) is given as
$PCV_{i}=N_{i}(\hat{\theta}_{0,i}+\bar{p}),$ (20)
where $N_{i}$ is the number of involvements of players $i$ in possession over the course of the season and $\bar{p}$ is the proportion of goals scored from the possessions. The rationale behind the PCV metric is the following: Considering a general possession, where no details are known about the possession, a good estimate for the probability of scoring is to simply take the proportion of scoring a goal from a possession in our dataset, so $\bar{p}$. Since $\hat{\theta}_{0,i}$ represents the level shift in the probability of scoring a goal from a possession when player $i$ is involved in the possession, then if it is known that player $i$ is involved in this general possession, adding the level shift $\hat{\theta}_{0,i}$ to $\bar{p}$ represents a more accurate estimate of the scoring probability. Multiplying this by the number of involvements results in an estimate for the average contribution to possessions by a player over the course of the observed period. Of course, a downside of such an adjustment is that players who played fewer matches during this period due to injuries, for example, might be undervalued. This effect can be alleviated by e.g. standardizing the involvements to a certain number of minutes played. Finally we note that the design matrix $X_{mod}$ in Bajons (2023) is slightly different to the one used in this paper.
A.2 Comparison of the PCV rating with the regularized regression models
Table 8 shows the top 5 players of each position group with regards to the $PCV$ value from Equation (20). As expected, forwards and midfielders obtain a better score than defenders and goalkeepers. A glance at Table 8 suggests that the results are similar to the ones from the ridge and the group lasso regression. However, Figure 5 provides a more in-depth comparison. First, it can be seen that the $PCV$ estimates are highest correlated to the direct involvement coefficients and that there seems to be very little relation between $PCV$ and the OF coefficients from the regression approaches. While this observation is in accordance with the intuition, the figure also demonstrates that there is little agreement between the regression estimates and DML ratings for defensive players. The main implication of these insights is that under the DML rating, it is more difficult to accurately evaluate the contributions of defensive players. In that sense, the regression approaches are more flexible. An advantage of the DML approach is however that the rating can be interpreted as average contribution to possessions by a player over the course of the observed period, which may be more intuitive for practitioners than the coefficients from penalized regression models.
Table 8: Top 5 of each position group (forward, midfielder, defender, goalkeeper) under the PCV rating derived from the debiased machine learning model.
Position Player PCV Score Rank
Forward L. Messi 43.23 1
L. Suárez 35.02 2
Iago Aspas 31.75 3
Cristiano Ronaldo 29.87 4
Maxi Gómez 28.30 5
Midfielder D. Wass 28.00 6
T. Kroos 27.51 7
Álex Granell 21.41 12
É. Banega 20.74 13
Samu Castillejo 20.37 14
Defender Jonny 21.61 11
C. Lenglet 15.28 38
Barragán 15.13 39
Álvaro González 14.85 41
Piqué 14.20 50
Goalkeeper M. ter Stegen 9.57 111
K. Navas 9.57 112
Adán 5.69 196
M. Dmitrović 5.07 218
J. Oblak 5.04 220
Figure 5: Correlation of the Ratings from the DML approach with the ones from ridge regression (x-axis) and the group lasso (y-axis). Each of the 4 position groups (forwards, midfielders, defenders, goalkeepers) are colored differently and each of the involvement types (Inv, OF, Sum) from the regression approaches is displayed by a different symbol.
Finally, as is observed in the main text (Section 4.3), the $PCV$ metric as derived from the DML approach performs worse than the penalization models in terms of validity. It seems that it serves mostly as a more interpretative metric, which is often preferred by practitioners. The correlation of the ratings of offensive players (forwards and midfielders) to ridge and group lasso regression indicates that at least for these groups the model provides meaningful results. In addition, the results in Bajons (2023) suggest that a different data setup may lead to improved results in terms of validity. Nevertheless, this approach is not able to match the flexibility in terms of interpretation that is gained from the regularized methods.
Appendix B Scouting players using the exclusive lasso penalty
As mentioned in the Section 3.2 the exclusive lasso offers a lot of potential for scouting players. In order to illustrate this, we consider the special case of grouping the players according to the team they play for, the possession group (as in the Section 4.2.1), and of course also the direct and indirect involvement. Since there are 20 teams playing in La Liga, this leads to $20\cdot 4\cdot 2=160$ player groups. Due to the structure of the exclusive lasso, such a grouping thus selects the most representative (i.e. the strongest resp. weakest) players for each team and on each position, all other players will obtain a coefficient of 0. For scouting this is appealing as the algorithm leaves only a range of players from each team that may be considered for recruitment. Note that teams interested in scouting players could group them in any kind of way that is of interest. Another example would be the case, where a team is interested in scouting a player playing a specific position, that is younger than a certain age and whose market value is below some budget bound. Then the team could easily form player groups accordingly while keeping other peer groups (such as team specific player position groups for example) for comparison.
Table 9: Top Players of 18 of the 20 teams of the spanish La Liga with respect to the sum of direct (Inv) and indirect (OF) involvement.
Score
Player Inv OF Sum Position Team
L. Messi 0.24604 0.15618 0.40222 Forward Barcelona
Cristiano Ronaldo 0.35760 0.00000 0.35760 Forward Real Madrid
A. Griezmann 0.26049 0.00000 0.26049 Forward Atlético Madrid
J. Murillo 0.00000 0.20968 0.20968 Defender Valencia
Juanmi 0.15745 0.03252 0.18996 Forward Real Sociedad
Iago Aspas 0.17601 0.00000 0.17601 Forward Celta de Vigo
Portu 0.17180 0.00000 0.17180 Midfielder Girona
C. Bacca 0.22732 -0.0613 0.16603 Forward Villarreal
Charles 0.15551 0.00000 0.15551 Forward Eibar
A. Sanabria 0.14040 0.00000 0.14040 Forward Real Betis
W. Ben Yedder 0.16682 -0.0401 0.12672 Forward Sevilla
Munir 0.23037 -0.10872 0.12165 Forward Deportivo Alavés
Ángel 0.11359 0.00000 0.11359 Forward Getafe
Rubén Pérez -0.03250 0.14109 0.10859 Midfielder Leganés
C. Borges 0.01045 0.08178 0.09222 Midfielder Deportivo La Coruña
E. Boateng 0.07414 0.00000 0.07414 Forward Levante
Granero 0.00000 0.06252 0.06252 Midfielder Espanyol
Iñaki Williams 0.05245 0.00000 0.05245 Forward Athletic Club
Table 9 displays the best player from each of the 20 teams of La Liga with respect to the sum of their on-ball and off-ball contributions [^5] Two teams (Málaga and Las Palmas) have only players with negative sum coefficients and thus are not depicted in this table.. From a domain-specific view, the results are reasonable and nicely interpretable. Most of the top players from the teams are forwards or midfielders. Furthermore, the best-ranked players are also the ones from the best teams that season in terms of final standings. Only from Valencia, the best player is a defender. Interestingly, Valencia played a very good season especially defensively, receiving relatively few goals (only 3 teams received fewer goals this season).
Table 10: Top 11 players selected with respect to direct involvement.
Score
Player Inv OF Sum Position Team
Cristiano Ronaldo $0.358$ $0.000$ $0.358$ Forward Real Madrid
Maxi Gómez $0.272$$0.149$ $0.124$ Forward Celta de Vigo
A. Griezmann $0.260$ $0.000$ $0.260$ Forward Atlético Madrid
L. Messi $0.246$ $0.156$ $0.402$ Forward Barcelona
Munir $0.230$$0.109$ $0.122$ Forward Deportivo Alavés
C. Bacca $0.227$$0.061$ $0.166$ Forward Villarreal
C. Stuani $0.219$$0.185$ $0.034$ Forward Girona
Lucas Pérez $0.212$$0.150$ $0.062$ Forward Deportivo La Coruña
Rodrigo $0.197$ $0.000$ $0.197$ Forward Valencia
L. Suárez $0.184$ $0.000$ $0.184$ Forward Barcelona
Iago Aspas $0.176$ $0.000$ $0.176$ Forward Celta de Vigo
When using the results for scouting or analyzing players there are however a few peculiarities to keep in mind. As described in the Section 3.2 the penalty ensures that on a group level, a ridge-like behavior is observed. That is, on the one hand, this means that in every group there is at least one representative which has a coefficient value different from zero. On the other hand, the group effect is shrunken towards zero, where more influential groups obtain less shrinkage. Within a group however, the coefficients are competitively set towards zero, such that only the most relevant group members obtain a coefficient value different from zero. For more intuition on the behavior of the exclusive lasso penalty, we refer to Campbell and Allen (2017). That being said, analyses or comparisons of individual coefficients, such as e.g. the direct involvement coefficients of specific players, have to be conducted with care. Table 10 displays the top 11 players with respect to the on-ball coefficients. Judging purely by coefficient, it seems that Cristiano Ronaldo is way stronger than his peers. However, a further look at the table shows that from two teams (Barcelona and Celta de Vigo) two forwards are in the top 11. The exclusive lasso penalty acts in a way such that the magnitude of the (shrunken) effect of a group is divided between all relevant representatives of said group. That is, if there is more than one player in a group that obtains a value different from zero, then the whole group effect is competitively shared by all of them.
Table 11: Top 5 player-team groups form the 80 direct involvement groups.
Group Effect # Group Members
Team Position Inv OF Inv OF
Celta de Vigo Forward 0.44839 -0.14855 2 1
Barcelona Forward 0.42982 0.12873 2 2
Real Madrid Forward 0.35760 0.13353 1 1
Atlético Madrid Forward 0.26049 0.10014 1 1
Deportivo Alavés Forward 0.23037 -0.07741 1 2
Table 11 shows the player-team groups with the highest magnitude of effects with respect to the direct involvement. The two best groups (the forwards from Celta de Vigo and the forwards from Barcelona) both have two selected players, whereas, for the other groups, the algorithm chose only one representative. For the two groups, the players that have coefficients different from zero compete for the magnitude of the group effect and thus it is not straightforward to compare their coefficients directly to the ones from groups with only one selected player. In order to have a fair comparison between e.g. the two selected forwards from Barcelona and the one selected forward from Real Madrid, these players would need to be merged into one group [^6] Compare e.g. with the results from Section 4.2.1..
The setting discussed in this appendix provides some intuition on the usage of the exclusive lasso. Due to the internal functionality of this penalty, there are a lot of interesting potential use cases, especially for scouting players. From the perspective of a practitioner, it is attractive that the algorithm selects only a few strong players from each group. This leads to quick identification of interesting players that may be analyzed in more detail by scouting experts and the player development staff. When trying to interpret the effects of the exclusive lasso penalty there are however pitfalls that one needs to be aware of.

### v2

@@ 5 Discussion
In this paper, we present a novel framework for evaluating the strength of football players. We adapt the ideas of classical regularized adjusted plus-minus models to possession sequences data and employ a diverse set of penalization structures. The former adjustment allows for evaluating on-ball as well as off-ball contributions of players – an often overlooked but important part of the game. The latter allows for more flexibility and the incorporation of domain-specific characteristics of the game by taking into account e.g.~positions of players.
In practice, each penalization structure has its advantages respectively use cases. The general picture is, however, similar throughout all presented models. The on-ball contributions are dominated by offensive players (forwards and midfielders), while defensive players score better on the off-ball contributions. From a domain-specific viewpoint, this result is expected as the measurement for evaluating possession is whether they result in a goal or not. It is reassuring that all the methods act similarly in that sense and these results suggest that they are meaningful and informative. Furthermore, we provide objective justification for using our framework by conducting a statistically-informed tailored-to-soccer validity analysis. The results thereof suggest that using the group information of players provides value and should be accounted for. Furthermore, the additional information of on-ball and off-ball value can be leveraged in several ways. For example, defensive players who perform well in on-ball contributions may identify as offensively oriented defenders, which might be desirable for specific styles of play. Similarly, offensive players that demonstrate high off-ball value may have more impact on the game than their active contributions (as measured through goals, assists, etc.) are disclosing.
@@
The flexibility of our approach combined with the ease of adapting it to existing procedures yields several possible future directions to boost football player performance evaluation. For one, it might be interesting to simply adapt existing regularized adjusted PM models to allow for the presented regularization schemes. Also incorporating adjustments for home advantage, manpower, league strength or relevancy of performance (cf. Kharrat et al., 2020; Pantuso and Hvattum, 2021) is straightforward. However, this becomes more relevant when more data is available. Another interesting direction is to adapt the target variable and allow for some sort of hybrid approach. One possibility is to follow Kharrat et al. (2020) and use xG values as target variable, however, any bottom-up rating that assigns a value at the end of the possession may be used. One could even think of ordinal-type target variables, however in this case the application of the more elaborated penalization structures needs to be developed further. Adapting the target may help discriminate between players, as considering goals as outcome variables leads to a sparse and imbalanced setup. Similarly, since analyzing possessions provides more detailed insights into the game than analyzing segments as defined in PM models, it is possible to adjust the outcome variable to allow for the valuation of defensive actions such as the ability of a team to stop dangerous possessions. Finally, future endeavors may explore and compare different setups for incorporating player information. For example, it might be an interesting avenue to account for the off-ball involvement by separating offensive and defensive participation on the field (indirect involvement). Figuratively speaking, a player like Lionel Messi, while without a doubt positively influencing an offensive possession simply by being a threat on the field, has arguably only little effect with regard to defensive off-ball contribution.
@@
In summary, we believe that the presented framework allows for novel perspectives and enhancements in the field of player evaluation and scouting. The results are not only from a domain-specific but also from an objective perspective useful, and thus the stakeholders involved in the sports of football such as clubs, betting companies, media as well as fans may benefit from the presented methods.